A hospital received a shippment of 8-milligram doses of medicine. Each 8-milligram package was repacked into two smaller dosesof unequal size and labled packet a and packet b. The hospital then used 17 doses of packet A and 14 doses of packet B in one week. The hospital used a total of 127 milligrams of the medicine during that week. How many milligrams of the medicine are contained in each A packet? In each B packet?
Answer by stanbon(75887)

A hospital received a shipment of 8-milligram doses of medicine. Each 8-milligram package was repacked into two smaller doses of unequal size and labled packet a and packet b. The hospital then used 17 doses of packet A and 14 doses of packet B in one week. The hospital used a total of 127 milligrams of the medicine during that week. How many milligrams of the medicine are contained in each A packet? In each B packet
---
a + b = 8
17a + 14b = 127
------
Modify:
17a + 17b = 17*8
17a + 14b = 127
-----
Subtract and solve for "b":
3b = 9
b = 3
-----
Solve for "a":
a+b = 8
a + 3 = 8
a = 5
-----------
Ans: 3 "a" packs and 5 "b" packs.
=====================================
Cheers,
Stan H.
========